OpenStudy (anonymous):

can someone do this for me please? but show the steps, too please~ ^-^ Find the solution to the system by LINEAR COMBINATION: 4x + 3y = (-2) 3x + 2y = (-3)

OpenStudy (anonymous):

4x + 3y = (-2) [1] 3x + 2y = (-3) [2] Okay let's start by getting rid of the x variable 3*[1], 12x + 9y = (-6) [3] 4*[2], 12x + 8y = (-12) [4] Now we are going [3]-[4] 9y-8y=-6-(-12) y=-6+12 =6 Then sub y=6 back into either [1] or [2] 3x+2*6=-3 3x+12=-3 3x=-15 x=-5 So x=-5 and y=6

OpenStudy (anonymous):

ty~ I was having troubles with this one for some reason :D